# Unifying Autoregressive and Diffusion-Based Sequence Generation

### v1

@@ 3.1 Hyperschedules
3.2 Hybrid Processes
In the absorb process, each step produces the curriculum by replacing some of $\mathbf{y}$’s entries by the special MASK token. Conversely, at generation time, the only action available to the generator is to replace some MASK tokens by non-MASK ones. Notice that, unless the generator is “perfect”, it may become apparent late in the generation process that some early token choices were, in retrospect, inherently incompatible. However, there are no action available for the model to “fix” these token choices: no backsies. Although such “hindsight” situations may occur in any domain, they are particularly relevant to computer code generation and other reasoning-intensive tasks. [^3] As an extreme example, consider the graph coloring of a particularly nasty instance.
@@ 3.2 Hybrid Processes
These observations motivate an hybrid forward process, of which we consider two varieties. The first one uses the SEDD framework with $Q_{\text{tok}}$ set to $Q_{\text{Hybrid}}^{\gamma{}}=(1-\gamma)Q_{\text{Absorb}}+\gamma Q_{\text{Uniform}}$, interpolating the $Q_{\text{Uniform}}$ and $Q_{\text{Absorb}}$ extremes according to an hyperparameter $10<\gamma<$ 1. The evolution operator $\exp\bigl{(}\bigl{(}\bar{\sigma}_{\tau_{t}^{i}}-\bar{\sigma}_{\tau_{t+1}^{i}}\bigr{)}Q_{\text{Hybrid}}^{\gamma{}}\bigr{)}$ can be solved analytically (because $Q_{\text{Uniform}}$ and $Q_{\text{Absorb}}$ commute, see Appendix A), enabling use in practice with the standard SEDD loss.
Our second hybrid process variety is closer to the MDM framework: unless $\tau_{t}^{i}=0$ (in which case $x_{t}^{i}=y^{i}$), the probability distribution for $x_{t}^{i}$ is given by the $y^{i}$-th column of $\bigl{(}\mathds{1}+\epsilon Q_{\text{Uniform}}\bigr{)}\bigl{(}\mathds{1}+(1-\alpha\raisebox{2.0pt}{${}_{\tau_{t}^{i}}$})Q_{\text{Absorb}}\bigr{)}$, where $10<\epsilon<$ 1 is a step-independent probability that the token is substituted by a uniform one, followed by a standard MDM process henceforth. A weighted cross-entropy loss is used (see Appendix B.2) and, like MDMs, the neural network does not require an explicit noise-level dependency. Which of the two variety is used can be inferred from which of $\gamma$ or $\epsilon$ is specified.
@@ 3.3 Attention Mask and Efficiency
However, new opportunities for optimization come up when the hyperschedule follows a certain autoregressive-like regular structure such as the ones seen in Fig. 2(d)-2(c). More specifically, at each steps $t$ the hyperschedule $\boldsymbol{\tau}_{t}$ and the state $\hat{\mathbf{x}}_{t}$ both break in three components
@@ 4.2.1 Ablation Study
(a) Baseline SEDD-Absorb [12] 24.10 00 $\boldsymbol{\tau}_{\text{Flat}}$
[…]
@@ 4.2.2 Language Modeling Performance
Diffusion (Ours) $\gamma$-Hybrid [$\gamma{=0.02},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] (56B tokens) 110M 27.8
$\gamma$-Hybrid [$\gamma{=0.02},\boldsymbol{\tau}_{\text{Flat}},\textsc{shifted}$] (56B tokens) 110 M 28.3
$\gamma$-Hybrid [$\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{aligned}$] (65B) 110M 27.1
$\gamma$-Hybrid [$\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{aligned}$] (65B) 110M 27.0
$\gamma$-Hybrid [$\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{shifted}$] (65B) 110M 27.5
$\gamma$-Hybrid [$\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{shifted}$] (65B) 110M 26.6
@@ 4.3 Zero-Shot Generalization
[…]
@@
[…]
@@ 4.4 Sampler Performance
[…]
@@
$\epsilon$-Hybrid [$\epsilon=0.01$, $\boldsymbol{\tau}_{\text{Flat}}$, Aligned] Original Sampler 0.848 0.779 121.90 129.52 5.49 5.50
@@
$\epsilon$-Hybrid [$\epsilon=0.01$, $\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{4}}}}$, Aligned] Original Sampler 0.778 0.847 139.64 142.13 5.43 5.46
@@
[…]
@@ 4.5 Sequence Generation Quality-Diversity Trade-offs
We further investigate the trade-off between sequence generation quality and diversity by analyzing generative perplexity against token-level entropy and MAUVE scores (Pillutla et al., 2021), as visualized in Fig. 4. Our hybrid configurations consistently achieve superior positions on both Pareto frontiers, indicating enhanced generation fluency, coherence, and diversity compared to baselines. The observed balance in generation quality and diversity establish a new state-of-the-art in diffusion-based sequence generation.
@@ Appendix A Evolution Operators
$\operatorname{e}^{\Delta Q_{\text{Hybrid}}^{\gamma{}}}$ (7a)
@@ Appendix B Implementation Nuances
Figure 5: Two transformer-based sequence generators for $d=4$. (5(a)) The Aligned configuration of standard diffusion models is reminiscent of masked language models. (5(b)) The Shifted configuration is closer to autoregressive language models. Here $\hat{x}^{-1}$ represent a token solely part of the conditioning (i.e., not generated), and may or may not be constant (e.g., BOS). Similarly, Refer to caption represents that the output associated with the last token is discarded. Our position-based indexing abstracts away these details.
@@ B.1 Training Setup
We now detail our experimental training procedure, structured into two distinct stages. In Stage 1, we initially train our base models in both aligned and shifted configurations using the hybrid noising process paired with the flat hyperschedule $\tau_{\mathrm{flat}}$. Specifically, we adopt different modeling strategies depending on the chosen variant. For $\gamma$-variant models, we extend the existing discrete diffusion framework from Lou et al. (2024), integrating our proposed hybrid transition operator $Q_{\text{Hybrid}}^{\gamma{}}$. For $\epsilon$-variant models, we instead employ the newly proposed hybrid diffusion cross-entropy (HDCE) loss, detailed in Equation 8. Stage 1 models are trained separately on two standard datasets: OpenWebText and LM1B. Each configuration undergoes training for approximately 850K gradient updates with a batch size of 512, processing roughly  [^4] B tokens for OpenWebText and 56B tokens for LM1B.
In Stage 2, we fine-tune the Stage 1 models under alternative hyperschedules—specifically, $\tau_{\mathrm{block}}$ and $\tau_{\mathrm{slide}}$. These experiments involve custom-designed attention masks tailored to each hyperschedule. Due to this customization, highly optimized attention kernels such as Flash-Attention are not applicable, necessitating reliance on the standard PyTorch attention mechanism, which incurs higher computational costs. Stage 2 training continues for an additional 150K gradient steps, resulting in a cumulative training volume of 524B tokens for OpenWebText and 65B tokens for LM1B.
@@ B.2 Loss Function
$w_{i,t}=\begin{cases}\displaystyle\frac{1}{p_{\text{mask}}(i,t)}&\text{if }x_{i,t}\text{ is masked},\\[2.84526pt] \displaystyle\frac{\lambda(1-\epsilon)}{1-p_{\text{mask}}(i,t)}&\text{if }x_{i,t}\text{ is unmasked and shuffled},\\[2.84526pt] \displaystyle\frac{\lambda\epsilon}{1-p_{\text{mask}}(i,t)}&\text{if }x_{i,t}\text{ is unmasked and not shuffled},\end{cases}$ (9)
@@ B.3 Efficient Training and Inference
Figures 9 and 10 present the corresponding matrices for the Shifted configuration. Notice how settled tokens (light red or the gray $\hat{x}^{-1}$) are repeated as the first input of an interval to denoise in the second half of the matrix, and how the last output of each such interval is discarded.
B.4 Inference and KV-Caching
At inference, both Euler and $\tau$-leaping analytical solutions are available; however, our empirical results suggest that $\tau$-leaping is the de facto superior choice. As a result of the presence of $\hat{\mathbf{x}}_{t}^{\text{settled}}$ tokens, we can leverage KV-caching to accelerate inference. Specifically, during each forward pass of the transformer, only the $\hat{\mathbf{x}}_{t}^{\text{active}}$ tokens are updated while the cached keys and values for $\hat{\mathbf{x}}_{t}^{\text{settled}}$ remain unchanged.
@@ C.1 Upper Bound Estimation of Perplexity
10:     $\mathcal{L}_{\text{batch}}\leftarrow\frac{1}{M}\,\mathcal{L}_{\text{batch}}$triangleright Average over MC samples
@@ F.1 Zeroshot Perplexity
[…]

### v2

@@ 3.1 Hyperschedules: position-dependent schedules
3.2 Hybrid Processes: training denoisers that can fix their mistakes
In the absorb process, each step produces the curriculum by replacing some of $\mathbf{x}_{T}$’s entries by the special MASK token. Conversely, at generation time, the only action available to the generator is to replace some MASK tokens by non-MASK ones. Notice that, unless the generator is “perfect”, it may become apparent late in the generation process that some early token choices were, in retrospect, inherently incompatible. However, there are no action available for the model to “fix” these token choices: no backsies. Although such “hindsight” situations may occur in any domain, they are particularly relevant to computer code generation and other reasoning-intensive tasks. [^1] As an extreme example, consider the graph coloring of a particularly nasty instance.
@@ 3.2 Hybrid Processes: training denoisers that can fix their mistakes
These observations motivate an hybrid forward process, of which we consider two varieties. The first one uses the SEDD framework with $Q_{\text{tok}}$ set to $Q_{\text{Hybrid}}^{\gamma{}}=(1-\gamma)Q_{\text{Absorb}}+\gamma Q_{\text{Uniform}}$, interpolating the $Q_{\text{Uniform}}$ and $Q_{\text{Absorb}}$ extremes according to an hyperparameter $0<\gamma<1$. The evolution operator $\exp\bigl(\bigl(\bar{\sigma}_{\tau_{t}^{i}}-\bar{\sigma}_{\tau_{t+1}^{i}}\bigr)Q_{\text{Hybrid}}^{\gamma{}}\bigr)$ can be solved analytically (because $Q_{\text{Uniform}}$ and $Q_{\text{Absorb}}$ commute, see Appendix A), enabling use in practice with the standard SEDD loss.
Our second hybrid process variety is closer to the MDM framework: unless $\tau_{t}^{i}=0$ (in which case $x_{t}^{i}=x_{T}^{i}$), the probability distribution for $x_{t}^{i}$ is given by the $x_{T}^{i}$-th column of $\bigl(\mathds{1}+\epsilon Q_{\text{Uniform}}\bigr)\bigl(\mathds{1}+(1-\alpha\raisebox{2.0pt}{${}_{\tau_{t}^{i}}$})Q_{\text{Absorb}}\bigr)$, where $0<\epsilon<1$ is a step-independent probability that the token is substituted by a uniform one, followed by a standard MDM process henceforth. A weighted cross-entropy loss is used (see Appendix B.2) and, like MDMs, the neural network does not require an explicit noise-level dependency. Which of the two variety is used can be inferred from which of $\gamma$ or $\epsilon$ is specified.
@@ 3.3 Attention Mask and Efficiency
However, new opportunities for optimization come up when the hyperschedule follows a certain autoregressive-like regular structure such as the ones seen in Figs. 2(c.1, d.1). Indeed, at each steps $t$ the hyperschedule $\boldsymbol{\tau}_{t}$ and the state $\hat{\mathbf{x}}_{t}$ both break in three components
@@ 4.2.1 Ablation Study
(a) Baseline SEDD-Absorb [14] 24.10 0 $\boldsymbol{\tau}_{\text{Flat}}$
(b) (a) $\mathbf{+}$ Hybrid Process 22.30 $0.01$ $\boldsymbol{\tau}_{\text{Flat}}$
(c) (b) $\mathbf{+}$ Weighted token-embedding(= $\gamma$-Hybrid) 22.18 $0.01$ $\boldsymbol{\tau}_{\text{Flat}}$
(d) (c) $\mathbf{-}$ Transformer time-conditioning 22.47 $0.01$ $\boldsymbol{\tau}_{\text{Flat}}$
(e) (c) $\mathbf{+}$ $\boldsymbol{\tau}_{\text{Slide}}^{\omega{=d}}$ 21.53 $0.01$ $\boldsymbol{\tau}_{\text{Slide}}^{\omega{=d}}$
@@ 4.2.2 Language Modeling Performance
Diffusion (Ours) $\gamma$-Hybrid $\gamma{=0.02},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] (56B tokens) 110M 27.8
$\gamma$-Hybrid $\gamma{=0.02},\boldsymbol{\tau}_{\text{Flat}},\textsc{shifted}$] (56B tokens) 110 M 28.3
$\gamma$-Hybrid $\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{aligned}$] (65B) 110M 27.1
$\gamma$-Hybrid $\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{aligned}$] (65B) 110M 27.0
$\gamma$-Hybrid $\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{shifted}$] (65B) 110M 27.5
$\gamma$-Hybrid $\gamma{=0.02},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{shifted}$] (65B) 110M 26.6
@@ 4.3 Zero-Shot Generalization
We evaluate the generalization capabilities of our models across seven standard datasets considered before by Lou et al. (2024) and Sahoo et al. (2024). As shown in Table 3, our Hybrid diffusion language family of models, ($\gamma\text{-Hybrid}$ and $\epsilon\text{-Hybrid}$) consistently surpasses prior discrete diffusion approaches on 5 of the 7 benchmarks. Moreover, we narrow the gap between diffusion-based and AR models, outperforming AR baselines on two datasets.
@@
Transformer      (Sahoo et al., 2024) 82.05 25.75 51.25 51.28 52.09 49.01 41.73
SEDD Absorb        (Lou et al., 2024) 96.33 35.98 68.14 48.93 67.82 45.39 40.03
MDLM     (Sahoo et al., 2024) 90.96 33.22 64.94 48.29 62.78 43.13 37.89
BD3-LM $L^{\prime}=4$        (Arriola et al., 2025) 96.81 31.31 60.88 50.03 61.67 42.52 39.20
$\gamma$-Hybrid ( [^4] B)        $\gamma{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] 89.94 30.02 61.01 45.38 67.51 46.57 40.62
$\epsilon$-Hybrid ( [^4] B)        $\epsilon{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] 90.89 32.53 68.91 50.23 64.61 41.18 37.85
$\gamma$-Hybrid     $\gamma{=0.01},\boldsymbol{\tau}_{\text{Slide}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{aligned}$] 90.67 31.73 73.71 50.03 68.27 41.49 37.89
$\gamma$-Hybrid     $\gamma{=0.01},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{shifted}$] 95.22 32.64 63.68 44.75 62.18 42.01 37.33
@@ 4.4 Sampler Performance
We evaluate the performance gained from our proposed adaptive correction sampler (ACS). Table 4 empirically shows that ACS improves the $\epsilon\text{-Hybrid}$ models’ ability to correct initial missteps, resulting in sequences that are both more coherent and of higher quality. Pseudocode for the original sampler and ACS can be found in Appendix D.
@@
$\epsilon$-Hybrid $\epsilon=0.01$, $\boldsymbol{\tau}_{\text{Flat}}$, Aligned] Original Sampler 0.848 0.779 121.90 129.52 5.49 5.50
@@
$\epsilon$-Hybrid $\epsilon=0.01$, $\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{4}}}}$, Aligned] Original Sampler 0.778 0.847 139.64 142.13 5.43 5.46
@@
Table 4: Comparing the original sampler with ACS on $\epsilon\text{-Hybrid}$ variants. For a more complete comparison please refer to Table 10.
@@ 4.5 Sequence Generation Quality-Diversity Trade-offs
We further investigate the trade-off between sequence generation quality and diversity by analyzing generative perplexity against token-level entropy and MAUVE [^2] MAUVE strongly correlates with human judgment of text quality and diversity. scores (Pillutla et al., 2021), as visualized in Fig. 3. Our hybrid configurations consistently achieve superior positions on both Pareto frontiers, indicating enhanced generation fluency, coherence, and diversity compared to baselines; a new state-of-the-art in diffusion language modeling.
@@ Appendix A Evolution Operators
$\operatorname{e}^{\Delta Q_{\text{Hybrid}}^{\gamma{}}}\hskip-14.22636pt$ (7a)
@@
Additional notes on ELBO and hyperschedules.
Theorem 3.6 from Lou et al. (2024) provides an ELBO bound enabling training with the DWDSE training loss. This Theorem holds for arbitrary constant-coefficients diffusion matrix, but is derived for the special case of position-independent schedules. One might thus wonder if the result still holds if we simply substitute in our position-dependent hyperschedules.
As mentioned in the beginning of Section 3.3 of Lou et al. (2024), such a general diffusion matrix would be exponential in size, justifying decomposing the process into $d$ independent subspaces in practical applications. The evolution operator for each of these subspaces has the form $\exp(\Delta Q_{\text{tok}})$ and, because of their independence, the full evolution operator is simply the product of their action on the independent subspaces.
When substituting in an hyperschedule instead of a schedule in SEDD’s framework, the only difference is that the evolution operator for each subspace, $\exp(\Delta_{i}Q_{\text{tok}})$, gains a dependency on the position $i$ associated with this subspace. Crucially, the subspaces are still independent, and the overall evolution operator is still a simple product of the independent solutions. All other derivations leading to SEDD’s ELBO bound thus directly generalize to hyperschedules, including the case of $\gamma$-Hybrid.
The case of $\epsilon$-Hybrid is more complicated. When $\epsilon=0$, we fall back on MDLM, and the ELBO derivations from Sahoo et al. (2024) thus apply. However, for $\epsilon>0$, the evolution operator cannot be represented under the form $\exp(\Delta Q_{\text{tok}})$: we do not have an ELBO bound for $\epsilon$-Hybrid when $\epsilon>0$. In any case, our empirical observations support its use in practice.
It would be interesting to see if training with a proper ELBO-bounded loss could improve the performance of $\epsilon$-Hybrid models, and if something akin to our ACS algorithm could be derived from first principles. However, these concerns are outside the scope of the present work.
@@ Appendix B Implementation Nuances
Figure 4: Two transformer-based sequence generators for $d=4$. (4(a)) The Aligned configuration of standard diffusion models is reminiscent of masked language models. (4(b)) The Shifted configuration is closer to autoregressive language models. Here $\hat{x}^{-1}$ represent a token solely part of the conditioning (i.e., not generated), and may or may not be constant (e.g., BOS). Similarly, Refer to caption represents that the output associated with the last token is discarded. Our position-based indexing abstracts away these details.
@@ B.1 Training Setup
We now detail our experimental training procedure, structured into two distinct stages. In Stage 1, we initially train our base models in both aligned and shifted configurations using the hybrid noising process paired with the flat hyperschedule $\tau_{\mathrm{flat}}$. Specifically, we adopt different modeling strategies depending on the chosen variant. For $\gamma$-variant models, we extend the existing discrete diffusion framework from Lou et al. (2024), integrating our proposed hybrid transition operator $Q_{\text{Hybrid}}^{\gamma{}}$. For $\epsilon$-variant models, we instead employ the newly proposed hybrid diffusion cross-entropy (HDCE) loss, detailed in Equation 8. Stage 1 models are trained separately on two standard datasets: OpenWebText and LM1B. Each configuration undergoes training for approximately 850K gradient updates with a batch size of 512, processing roughly  [^4] B tokens for OpenWebText and 56B tokens for LM1B. This pretraining took about 14 days on 8×A100-80 GB GPUs for OpenWebText and 7 days on 2×A100-80 GB GPUs for LM1B.
In Stage 2, we fine-tune the Stage 1 models under alternative hyperschedules—specifically, $\tau_{\mathrm{block}}$ and $\tau_{\mathrm{slide}}$. These experiments involve custom-designed attention masks tailored to each hyperschedule. Due to this customization, highly optimized attention kernels such as Flash-Attention are not applicable, necessitating reliance on the standard PyTorch attention mechanism, which incurs higher computational costs. Stage 2 training continues for an additional 150K gradient steps, resulting in a cumulative training volume of 524B tokens for OpenWebText and 65B tokens for LM1B. This fine-tuning required approximately 3 days on 8×A100-80 GB GPUs for OpenWebText and about 36 hours on 2×A100-80 GB GPUs for LM1B.
@@ B.2 Loss Function
[…]
@@ B.3 Efficient Training and Inference
Figures 8 and 9 present the corresponding matrices for the Shifted configuration. Notice how settled tokens (light red or the gray $\hat{x}^{-1}$) are repeated as the first input of an interval to denoise in the second half of the matrix, and how the last output of each such interval is discarded.
B.4 Inference and KV‐Caching
At inference, both Euler and $\tau$-leaping analytical solutions are available; however, our empirical results suggest that $\tau$-leaping is the de facto superior choice. Because each step “settles” a prefix of tokens, we can cache the corresponding key/value pairs and avoid recomputing them on subsequent passes. Concretely, let
$L=\text{sequence length},\quad\rho=\text{tokens generated per step},\quad\omega=\text{window width}.$
Then the number of transformer calls is
$N\;=\;\Bigl\lceil(L-\omega)/\rho\Bigr\rceil+1,$
and the total token‐processing cost is
Without cache: $N\times\omega,$
With KV‐cache: $\omega+(N-1)\,\rho.$
@@ C.1 Upper Bound Estimation of Perplexity
(To simplify the notation, the following uses $\mathbf{y}$ to represent non-noised samples.)
@@
10:  $\mathcal{L}_{\text{batch}}\leftarrow\frac{1}{M}\,\mathcal{L}_{\text{batch}}$ $\triangleright$ Average over MC samples
@@
Remarks 2.
All our “Test PPL” metrics are computed using the loss presented above, i.e., not the one presented in Appendix B.2. For this reason, the upper bound provided by Algorithm 1 assesses the quality of the neural network at language modeling. In particular, it does not account for the full inference strategy. Therefore, although the Test PPL metric may indirectly be affected by the training objective (described in Appendix B.2), the choice of hyperschedule and the values of $\epsilon$ or $\gamma$ through their combined impact on the training of the neural network’s weights, these hyperparameters are not directly involved when running Algorithm 1.
@@ F.1 Zeroshot Perplexity
GPT-2   (WebText) 138.43 41.60 75.20 45.04
Transformer      (Sahoo et al., 2024) 82.05 25.75 51.25 51.28 52.09 49.01 41.73
D3PM    (Austin et al., 2021) 200.82 75.16 138.92 93.47
Plaid    (Gulrajani & Hashimoto, 2024) 142.60 50.86 91.12 57.28
MD4   (Shi et al., 2024) 102.26 35.90 68.10 48.43
SEDD Absorb        (Lou et al., 2024) 96.33 35.98 68.14 48.93 67.82 45.39 40.03
MDLM     (Sahoo et al., 2024) 90.96 33.22 64.94 48.29 62.78 43.13 37.89
BD3-LM $L^{\prime}=4$        (Arriola et al., 2025) 96.81 31.31 60.88 50.03 61.67 42.52 39.20
RADD-$\lambda$-DCE        (Ou et al., 2024) 107.85 37.98 72.99 51.70
$\gamma$-Hybrid ( [^4] B)        $\gamma{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] 89.94 30.02 61.01 45.38 67.51 46.57 40.62
$\epsilon$-Hybrid ( [^4] B)        $\epsilon{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] 90.89 32.53 68.91 50.23 64.61 41.18 37.85
$\gamma$-Hybrid ( [^4] B)        $\gamma{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{shifted}$] 100.88 37.48 71.51 56.57 70.69 43.06 38.83
$\gamma$-Hybrid     $\gamma{=0.01},\boldsymbol{\tau}_{\text{Slide}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{aligned}$] 90.67 31.73 73.71 50.03 68.27 41.49 37.89
$\gamma$-Hybrid     $\gamma{=0.01},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{aligned}$] 95.32 38.94 70.49 48.18 67.32 44.23 42.78
$\gamma$-Hybrid     $\gamma{=0.01},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{aligned}$] 90.74 35.24 62.64 51.21 69.62 41.46 37.13
$\gamma$-Hybrid     $\gamma{=0.01},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{shifted}$] 95.22 32.64 63.68 44.75 62.18 42.01 37.33